OpenStudy (anonymous):

What is the length, in units, of segment CD? Answer choices: 6 8.5 7 9.8 @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

is there a picture?

OpenStudy (alexandervonhumboldt2):

5/7=7/x

OpenStudy (alexandervonhumboldt2):

9.8 think can you valculate by youself

OpenStudy (anonymous):

yeah I just cross multiply right?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

am i right

OpenStudy (anonymous):

Yes. I got 9.8 too. 5x=49 49/5 = 9.8 Thanks so much ! (:

OpenStudy (alexandervonhumboldt2):

http://openstudy.com/study#/updates/548b2de8e4b04ed7d6b44133